# Annealed Multiple Choice Learning: Overcoming limitations of Winner-takes-all with annealing

### v1

@@ 1 Introduction
Ambiguous prediction tasks arise in deep learning when the target $y$ is ill-defined from the input $x$. Predicting $y$ directly from $x$ can be challenging due to the partial predictability of $y$ from the information contained in $x$. Multiple Choice Learning (MCL) [24, 40] addresses these challenges by providing a small set of plausible hypotheses, each representing a different possible outcome given the input. MCL learns these hypotheses using a competitive training scheme that promotes the specialization of the hypotheses in distinct regions of the output space $\mathcal{Y}$. The framework iteratively partitions $\mathcal{Y}$ into a Voronoi tesselation and guides each hypothesis towards the barycenter of its respective Voronoi cell [60]. This mechanism makes MCL akin to a gradient-descent-based and conditional variant of the popular K-means algorithm [44]. Like K-means, MCL is sensitive to initialization, subject to hypothesis collapse [48], and more generally may converge towards arbitrarily suboptimal hypothesis configurations [60]. While there is a substantial body of literature addressing the limitations of K-means [3, 19, 55], relatively little has been done to address these challenges in the context of MCL [60, 48, 52]. The core issue of MCL lies with its greedy gradient-based update of the hypotheses. This greediness precludes the exploration of the hypothesis space, preventing MCL from optimally capturing the ambiguity of $y$. We propose to incorporate annealing into this gradient descent update in order to improve the robustness of MCL.
@@
Deterministic annealing [59] is a variant of simulated annealing. Annealing-based exploration relies on a sequence of random moves across the search space, whereas deterministic annealing seeks greater efficiency by replacing this random process with the exact minimization of a deterministic functional, namely the free energy of the system. It has been shown that deterministic annealing can be efficiently applied to clustering [58, 59]. In this article, we show that it can be further adapted to the conditional and gradient-based setting of MCL. The resulting algorithm, which we name aMCL for annealed MCL, addresses the main issues of MCL and achieves strong performance in practical settings while being straightforward to implement and amenable to analysis. Specifically, we make the following contributions.
@@
We provide extensive experimental validation, by applying this method i) to illustrative synthetic examples; ii) to a standard conditional quantization benchmark (the UCI datasets); and also iii) to the challenging audio task of speech separation (Section 5). The code will be made available.
@@ 2 Related Work
Multiple choice learning. MCL has been successfully applied to various machine learning tasks, typically using multi-head neural networks, with each head providing a prediction [39, 21, 41]. Several works observed the phenomenon of hypothesis collapse [8, 18, 32, 39, 63, 60], where some hypotheses are left unused during training. Various solutions have been proposed to tackle collapse [60, 48, 52]. Notably, [60] introduces Relaxed-WTA, which updates non-winning hypotheses with a gradient scaled by a small constant $\varepsilon$. However, this small gradient biases the hypotheses towards the global barycenter of the target distribution, which can be shown to be suboptimal [14].
@@ 3.2 Combining deterministic annealing with Multiple Choice Learning
Therefore, at the lowest level, aMCL simply consists of replacing the $\mathrm{min}$ operator from (2) by $\mathrm{softmin}$. aMCL introduces the temperature schedule as an additional hyperparameter. As highlighted by the literature on simulated annealing [25], it is crucial to ensure that temperature decreases slowly enough to benefit from the advantages of annealing. In practice, we obtained good results with linear and exponential schedulers (see also Section 5).
On a higher level, we can interpret the objective of aMCL as a smoothed version of the MCL objective. Smoothing with high temperature simplifies the optimization problem (6), reducing the number of local minima: we can conjecture from this analysis that aMCL will find a global minimum at high temperature, and we can expect it to stay optimal as long as the temperature decreases slowly enough [9]. We can also see aMCL as an input-dependent version of deterministic annealing [58, 59]. In this view, a high temperature encourages the exploration of the hypothesis space and mitigates the greediness of the gradient descent update (3). Moreover, following [25], we can posit that there exists an optimal temperature schedule striking a balance between exploration and optimization. Yet another interpretation is that aMCL constitutes an adaptative extension of Relaxed-MCL [60], where $\varepsilon=q_{T(t)}(f_{k}|x,y)$ depends both on the distance between the hypothesis $f_{k}(x)$ and the target $y$, and on the training step $t$. These interpretations shed light on the inner workings of aMCL. However, the complete training dynamic of the algorithm appears when we analyze aMCL through the lens of information theory and statistical physics, which is the purpose of the next Section.
@@ 4.1 Soft assignation and entropy constraints
Minimizing (1) is NP-hard [1, 10, 47]. Unsurprisingly, MCL can get trapped in local minima during training. In this Section, we discuss why the aMCL training scheme in Section 3.2 is more resilient to this pitfall. The first step towards our analysis is to observe that the $\mathrm{stop\_gradient}$ operator used in (6) of the aMCL update effectively turns the algorithm into an alternating optimization of the soft distortion
$D(q,f)=\int_{\mathcal{X}\times\mathcal{Y}}\sum_{k=1}^{n}\ell(f_{k}(x),y)\;q(f_{k}|x,y)\;p(x,y)\mathrm{d}x\mathrm{d}y\;,$ (7)
@@
Part $(i)$ states that the $\mathrm{softmin}$ operator is the solution of the entropy-constrained minimization of the soft distortion. Part $(ii)$ states that a necessary condition for minimizing the soft distortion is that each $f_{k}$ is a soft barycenter of the assignation distribution for each temperature $T$. Part $(iii)$ states that each gradient update moves $f_{k}$ towards this soft barycenter $f^{\star}$, and that the update speed depends on the probability mass $\gamma^{\star}_{k}$ of the points softly assigned to $f_{k}$. Together, they describe the training dynamics of aMCL. Note that as $T\rightarrow 0$, $q_{T}$ converges to a one-hot vector, and the soft barycenter in $(ii)$ becomes a hard barycenter. This is consistent with the necessary optimal condition for MCL, $f_{k}^{\star}(x)=\mathbb{E}_{Y\sim p(y\mid x)}[Y\;|\;Y\in\mathcal{Y}_{k}(x)]$, proved by Rupperecht et al. [60].
@@ 4.2 Rate-distortion curve
We have established in Section 4.1 that the aMCL training scheme is equivalent to an entropy-constrained alternating optimization of the soft-distortion (7), with each hypothesis $f_{k}$ moving towards a soft barycenter. In this Section, we describe the impact of temperature cooling on this training dynamic.
@@
Remarkably, this phenomenon occurs even at finite temperatures (see Appendix, Proposition 9). As the temperature decreases, a phenomenon of bifurcation occurs [58, 56]. During this process, the hypotheses iteratively split into sub-groups, as shown in Figure 4. The virtual number of hypotheses [57] for each $x$ is captured by the conditional rate-distortion function
$R_{x}(D^{\star})\triangleq\min_{\begin{subarray}{c}q\in\Delta_{n},f\in\mathcal{F}\\ D_{x}(q,f)\leq D^{\star}\end{subarray}}I(\hat{Y};Y)\;,$ (10)
where $D_{x}(q,f)=\int_{\mathcal{Y}}\sum_{k}\ell(f_{k}(x),y)q(f_{k}\mid x,y)p(y\mid x)\mathrm{d}y$, $I(\cdot,\cdot)$ is the mutual information, $Y\sim p(y\mid x)$ is the target distribution, and $\hat{Y}\sim q(f_{k}\mid x)$ denotes a distribution over $\mathcal{Y}$ of hypotheses location with $q(f_{k}\mid x)=\int_{\mathcal{Y}}q(f_{k}\mid x,y)p(y\mid x)\mathrm{d}y$. The rate-distortion function $R_{x}(D^{\star})$ has the following key properties.
@@
    $R_{x}$ is a non-increasing, convex and continuous function of $D^{\star}$, $R(D^{\star})=0$ for $D^{\star}\geq D_{x}^{\mathrm{max}}$, and for each $x$ the slope can be interpreted as $R^{\prime}_{x}(D^{\star})=-\frac{1}{T}$ when it is differentiable.
@@
The rate-distortion curve effectively describes the training trajectory of our algorithm, with the deterministic annealing procedure consisting of ascending along this curve [58, 56]. Interestingly, there is a set of critical temperatures at which the hypotheses suddenly split, increasing the number of sub-groups they form. In particular, the first of these critical temperatures corresponds to $D_{x}^{\mathrm{max}}$, a possible discontinuity point of the rate-distortion curve derivative (see Theorem 4.2.4 from [4]). By analogy with statistical physics, the behavior at these points has been coined phase transitions [22]. An illustration of the trajectory of MCL and aMCL in the rate-distortion space is shown in Figure 3. We see that MCL evolves along a constant rate $R=\log_{2}(n)$ (in bits) following $(a)$. In contrast, aMCL initially reaches the critical state at $D^{\star}=D_{x}^{\mathrm{max}}$ following $(b)$. After the transition, the trajectory of aMCL returns to the nominal rate following $(c)$. This seemingly more complex pattern hides a simpler optimization problem, which allows the distortion rate to vary.
These observations could be extended to the conditional settings by taking the average over the inputs $x\in\mathcal{X}$ of the conditional rate $R_{x}(D^{\star})$.
Figure 4: Conditional phase transitions with $t_{1}<t_{2}<t_{3}<t_{4}$. Results for a conditional version of the dataset in Figure 1, where the Gaussian moves linearly and increases with the input $x\in[0,1]$. aMCL was trained during $1000$ epochs, using a linear scheduler with $T_{\mathrm{ini}}=1.0$, and using $49$ hypotheses. Each subplot group corresponds to the predictions of the model at a given temperature, evaluated at different $x$ values. At temperature $T(t_{1})$, the hypotheses are at the barycenter. As temperature decreases they undergo a first phase transition (at temperature $T(t_{2})$ for $x=0.9$ and $T(t_{3})$ for $x=0.6$), moving toward each Gaussian’s barycenter, followed by a second phase transition at $T(t_{4})$. We see that earlier splits in the cooling schedule correspond to conditional distributions with higher variances.
@@ 4.3 Phase transitions
where $q(y\mid f_{k},x)$ denotes the posterior probability of assigning a point $y$ to the hypothesis $k$, calculated using Bayes’s rule [57]. At high temperatures, all hypotheses merge into the conditional barycenter of the distribution (9). In this setting, all the matrices $C_{k,k}$ are equal, $\mathscr{F}$ has a unique minimizer, and the stability of this global optimum is conditioned on the strict positivity of the Hessian of $\mathscr{F}$. The first critical temperature $T_{0}^{c}$ is defined as the first temperature for which the Hessian of $\mathscr{F}$ is no longer positive definite.
This temperature is connected with $D_{x}^{\max}$, the vanishing point of the rate-distortion function $R_{x}(D^{\star})$, which also corresponds to the optimal 1-hypothesis distortion. Indeed, $R_{x}(D^{\star})$ measures the virtual number of hypotheses, so that the first splitting of the hypotheses from a single point to many groups will coincide with the moment when $R_{x}(D^{\star})>0$. We summarize these observations in the following Proposition 4, which is illustrated in Figure 3.
@@
Noting $\lambda_{\max}(\cdot)$ the maximum eigenvalue of a matrix, we have:
$T_{0}^{c}(x)$ $\triangleq\inf\left\{T\mid\mathrm{inf}_{f\in\mathcal{F}}D_{x}(q_{T},f)\geq D_{x}^{\mathrm{max}}\right\}=2\lambda_{\mathrm{max}}(C_{1,1}(f,q\mid x))\;,$ (14)
$T_{0}^{c}$ $\triangleq\inf\left\{T\mid\mathrm{inf}_{f\in\mathcal{F}}D(q_{T},f)\geq\int_{x\in\mathcal{X}}D_{x}^{\mathrm{max}}p(x)\mathrm{d}x\right\}\leq 2\sup_{x\in\mathcal{X}}\lambda_{\mathrm{max}}(C_{1,1}(f,q\mid x))\;.$ (15)
After the first transition, an interesting phenomenon occurs for some distributions $p(x,y)$. Instead of splitting in all directions, the hypotheses $f_{k}$ continue to form a small number of subgroups. The hypotheses $f_{k}$ may undergo many phase transitions before they all split apart from each other: this is illustrated in Figure 4. Moreover, this recursive splitting reaches an end: under mild conditions [35], there is a critical temperature below which the hypotheses all separate from each other. This defines the last critical temperature $T_{\infty}^{c}$. Remarkably, $T_{\infty}^{c}$ is associated with the Shannon Lower Bound: the temperature at which $R_{x}(D^{\star})$ hits the lower bound $\mathrm{SLB}(D^{\star})$ corresponds to the moment when the hypotheses $f_{k}$ completely separate from each other (see Theorem 1 and 3 in [56]).

### v3

@@ 1 Introduction
Ambiguous prediction tasks arise in deep learning when the target $y$ is ill-defined from the input $x$. Predicting $y$ directly from $x$ can be challenging due to the partial predictability of $y$ from the information contained in $x$. Multiple Choice Learning (MCL) [25, 41] addresses these challenges by providing a small set of plausible hypotheses, each representing a different possible outcome given the input. MCL learns these hypotheses using a competitive training scheme that promotes the specialization of the hypotheses in distinct regions of the output space $\mathcal{Y}$. The framework iteratively partitions $\mathcal{Y}$ into a Voronoi tesselation and guides each hypothesis toward the barycenter of its respective Voronoi cell [63]. This mechanism makes MCL akin to a gradient-descent-based and conditional variant of the popular K-means algorithm [45]. Like K-means, MCL is sensitive to initialization, subject to hypothesis collapse [49], and more generally may converge toward arbitrarily suboptimal hypothesis configurations [63]. While there is a substantial body of literature addressing the limitations of K-means [3, 20, 58], relatively little has been done to address these challenges in the context of MCL [63, 49, 54]. The core issue of MCL lies with its greedy gradient-based update of the hypotheses. This greediness precludes the exploration of the hypothesis space, preventing MCL from optimally capturing the ambiguity of $y$. We propose to incorporate annealing into this gradient descent update in order to improve the robustness of MCL.
@@
Deterministic annealing [62] is a variant of simulated annealing. In simulated annealing, exploration relies on a sequence of random moves across the search space, whereas deterministic annealing seeks greater efficiency by replacing this random process with the exact minimization of a deterministic functional, namely the free energy of the system. It has been shown that deterministic annealing can be efficiently applied to clustering [61, 62]. In this article, we show that it can be further adapted to the conditional and gradient-based setting of MCL. The resulting algorithm, which we name aMCL for annealed MCL, addresses the main issues of MCL and achieves strong performance in practical settings while being straightforward to implement and amenable to analysis. Specifically, we make the following contributions.
@@
We provide extensive experimental validation, by applying this method i) to illustrative synthetic examples; ii) to a standard distribution estimation benchmark (UCI datasets); and also iii) to the challenging audio task of speech separation (Section 5). The accompanying code is made available. [^1] https://github.com/Victorletzelter/annealed_mcl
@@ 2 Related Work
Multiple choice learning. MCL has been successfully applied to various machine learning tasks, typically using multi-head neural networks, with each head providing a prediction [40, 22, 42]. Several works observed the phenomenon of hypothesis collapse [8, 19, 33, 40, 67, 63], where some hypotheses are left unused during training. Various solutions have been proposed to tackle collapse [63, 49, 54]. Notably, [63] introduces Relaxed-WTA, which updates non-winning hypotheses with a gradient scaled by a small constant $\varepsilon$. However, this small gradient biases the hypotheses toward the global barycenter of the target distribution, which can be shown to be suboptimal [14].
@@ 3.2 Combining deterministic annealing with Multiple Choice Learning
Therefore, at the lowest level, aMCL simply consists of replacing the $\mathrm{min}$ operator from (2) by $\mathrm{softmin}$. aMCL introduces the temperature schedule as an additional hyperparameter. As highlighted by the literature on simulated annealing [26], it is crucial to ensure that the temperature decreases slowly enough to benefit from the advantages of annealing. In practice, we experimented with both linear and exponential schedulers (see also Section 5).
On a higher level, we can interpret the objective of aMCL as a smoothed version of the MCL objective. Smoothing with high temperature simplifies the optimization problem (6), making the loss landscape easier to navigate: we can conjecture from this analysis that aMCL will find a global minimum at high temperature, and we can expect it to stay optimal as long as the temperature decreases slowly enough [9]. We can also see aMCL as an input-dependent version of deterministic annealing [61, 62]. In this view, a high temperature encourages the exploration of the hypothesis space and mitigates the greediness of the gradient descent update (3). Moreover, following [26], we can posit that there exists an optimal temperature schedule striking a balance between exploration and optimization. Yet another interpretation is that aMCL constitutes an adaptative extension of Relaxed-MCL [63], as $q_{T(t)}(f_{k}\;|\;x,y)$ depends both on the distance between the hypothesis $f_{k}(x)$ and the target $y$, and the training step $t$. These interpretations shed light on the inner workings of aMCL. However, the complete training dynamic of the algorithm appears when we analyze aMCL through the lens of information theory and statistical physics, which is the purpose of the next section.
@@ 4.1 Soft assignation and entropy constraints
Minimizing (1) is NP-hard [1, 10, 48]. Unsurprisingly, MCL can get trapped in local minima during training. In this Section, we discuss why the aMCL training scheme in Section 3.2 is more resilient to this pitfall. The first step toward our analysis is to observe that the $\mathrm{stop\_gradient}$ operator used in (6) of the aMCL update effectively turns the algorithm into an alternating optimization of the soft distortion
$D(q,f)\triangleq\int_{\mathcal{X}\times\mathcal{Y}}\sum_{k=1}^{n}\ell(f_{k}(x),y)\;q(f_{k}|x,y)\;p(x,y)\mathrm{d}x\mathrm{d}y\;,$ (7)
@@
Part $(i)$ states that the $\mathrm{softmin}$ operator is the solution of the entropy-constrained minimization of the soft distortion. Part $(ii)$ states that a necessary condition for minimizing the soft distortion is that each $f_{k}$ is a soft barycenter of the assignation distribution for each temperature $T$. Part $(iii)$ states that each gradient update moves $f_{k}$ toward this soft barycenter $f^{\star}$, and that the update speed depends on the probability mass $\gamma^{\star}_{k}$ of the points softly assigned to $f_{k}$. Together, they describe the training dynamics of aMCL. Note that as $T\rightarrow 0$, $q_{T}$ converges to a one-hot vector, and the soft barycenter in $(ii)$ becomes a hard barycenter. This is consistent with the necessary optimal condition for MCL, $f_{k}^{\star}(x)=\mathbb{E}_{Y\sim p(y\mid x)}[Y\;|\;Y\in\mathcal{Y}_{k}(x)]$, proved by Rupperecht et al. [63].
@@ 4.2 Rate-distortion curve
We have established in Section 4.1 that the aMCL training scheme is equivalent to an entropy-constrained alternating optimization of the soft-distortion (7), with each hypothesis $f_{k}$ moving toward a soft barycenter. In this Section, we describe the impact of temperature cooling on this training dynamic.
@@
Remarkably, this phenomenon occurs even at finite temperatures (see Appendix, Proposition 9). As the temperature decreases, a phenomenon of bifurcation occurs [61, 59]. During this process, the hypotheses iteratively split into sub-groups, as shown in Figure 4. The virtual number of hypotheses [60] for each $x$ at a given distortion level is captured by the conditional rate-distortion function
$R_{x}(D^{\star})\triangleq\min_{\begin{subarray}{c}q\in\Delta_{n},f\in\mathcal{F}\\ D_{x}(q,f)\leq D^{\star}\end{subarray}}I_{x}(\hat{Y};Y)\;.$ (10)
In Eq. (10), $Y\sim p(y\mid x)$ follows the target distribution, the hypothesis position $\hat{Y}\sim q(f_{k}\mid x)$ follows a distribution over $\mathcal{Y}$ with $q(f_{k}\mid x)=\int_{\mathcal{Y}}q(f_{k}\mid x,y)p(y\mid x)\mathrm{d}x$, $I_{x}(\hat{Y};Y)$ is their mutual information, and $D_{x}(q,f)=\int_{\mathcal{Y}}\sum_{k}\ell(f_{k}(x),y)q(f_{k}\mid x,y)p(y\mid x)\mathrm{d}y$ is the distortion for input $x$.
The rate-distortion function $R_{x}(D^{\star})$ has the following key properties.
@@
    $R_{x}$ is a non-increasing, convex and continuous function of $D^{\star}$, $R_{x}(D^{\star})=0$ for $D^{\star}\geq D_{x}^{\mathrm{max}}$, and for each $x$ the slope can be interpreted as $R^{\prime}_{x}(D^{\star})=-\frac{1}{T}$ when it is differentiable.
@@
The rate-distortion curve effectively describes the training trajectory of our algorithm, with the deterministic annealing procedure consisting of ascending along this curve [61, 59]. Interestingly, there is a set of critical temperatures at which the hypotheses suddenly split, increasing the number of sub-groups they form. By analogy with statistical physics, the behavior at these points has been coined phase transitions[23]. An illustration of the trajectory of MCL and aMCL in the rate-distortion space is shown in Figure 3. We see that MCL evolves along a constant rate $R=\log_{2}(n)$ (in bits) following $(a)$. In contrast, aMCL initially reaches the critical state at $D^{\star}=D_{x}^{\mathrm{max}}$ following $(b)$. After the transition, the trajectory of aMCL returns to the maximal rate following $(c)$. We expect the optimization at a lower rate to be simpler and this training trajectory to provide a better initialization for the set of hypotheses compared to the vanilla MCL.
Figure 4: Conditional phase transitions with $t_{1}<t_{2}<t_{3}<t_{4}$. Results for a conditional version of the dataset in Figure 1, where the Gaussian moves linearly and increases with the input $x\in[0,1]$. aMCL was trained during $1000$ epochs, using a linear scheduler with $T_{0}=1.0$, and using $49$ hypotheses. Each subplot group corresponds to the predictions of the model at a given temperature, evaluated at different $x$ values. At temperature $T(t_{1})$, the hypotheses are at the barycenter. As temperature decreases they undergo a first phase transition (at temperature $T(t_{2})$ for $x=0.9$ and $T(t_{3})$ for $x=0.6$), moving toward each Gaussian’s barycenter, followed by a second phase transition at $T(t_{4})$. We see that earlier splits in the cooling schedule correspond to conditional distributions with higher variances.
@@ 4.3 Phase transitions
where $q(y\mid f_{k},x)$ denotes the posterior probability of assigning a point $y$ to the hypothesis $k$, calculated using Bayes’s rule [60]. At high temperatures, all hypotheses merge into the conditional barycenter of the distribution (9). In this setting, all the matrices $C_{k,k}$ are equal to the data covariance matrix $C(x)\triangleq\mathbb{C}\mathrm{ov}_{(X,Y)\sim p(x,y)}[Y\mid X=x]$, $\mathscr{F}$ has a unique minimizer, and the stability of this global optimum is conditioned on the strict positivity of the Hessian of $\mathscr{F}$. The first critical temperature $T_{0}^{c}$ is defined as the first temperature for which the Hessian of $\mathscr{F}$ is no longer positive definite.
This temperature is connected to $D_{x}^{\max}$, the vanishing point of the rate-distortion function $R_{x}(D^{\star})$, which also corresponds to the optimal 1-hypothesis distortion [4]. Indeed, $R_{x}(D^{\star})$ measures the virtual number of hypotheses, so that the first splitting of the hypotheses from a single point to several groups will coincide with the moment when $R_{x}(D^{\star})>0$. We summarize these observations in the following Proposition 4, which is illustrated in Figure 3.
@@
See Definition 1, Propositions 6 $(ii)$ and 8 in Appendix. Let $\lambda_{\max}(\cdot)$ denote the maximum eigenvalue of a matrix, $D_{x}^{*}(T)\triangleq\mathrm{inf}_{f\in\mathcal{F}}D_{x}(q_{T},f)$, $D^{*}(T)\triangleq\mathrm{inf}_{f\in\mathcal{F}}D(q_{T},f)$, and $D_{\mathrm{max}}\triangleq\int_{x\in\mathcal{X}}D_{x}^{\mathrm{max}}p(x)\mathrm{d}x$. Then the two following properties hold.
  • (i)
    $D_{x}^{*}$ and $D^{*}$ are non-decreasing functions of $T$ and admit generalized inverses (with the convention $g^{-1}(\beta)=\mathrm{inf}\{\alpha\mid g(\alpha)\geq\beta\}$ for the generalized inverse of a function $g$).
  • (ii)
    The conditional (resp. non-conditional) first critical temperature $T_{0}^{c}(x)$ (resp. $T_{0}^{c}$) satisfy
    $T_{0}^{c}(x)$ $\triangleq(D_{x}^{*})^{-1}(D_{x}^{\mathrm{max}})=2\lambda_{\mathrm{max}}(C(x))\;,$ (14)
    $T_{0}^{c}$ $\triangleq(D^{*})^{-1}\left(D_{\mathrm{max}}\right)\leq 2\sup_{x\in\mathcal{X}}\lambda_{\mathrm{max}}(C(x))\;.$ (15)
After the first transition, an interesting phenomenon occurs for some distributions $p(x,y)$. Instead of splitting in all directions, the hypotheses $f_{k}$ continue to form a small number of subgroups. The hypotheses $f_{k}$ may undergo many phase transitions before they all split apart from each other: this is illustrated in Figure 4. Generally, this recursive splitting reaches an end: under mild conditions [36], there is a critical temperature below which the hypotheses all separate from each other. This defines the last critical temperature $T_{\infty}^{c}$. Remarkably, $T_{\infty}^{c}$ is associated with the Shannon Lower Bound: the temperature at which $R_{x}(D^{\star})$ hits the lower bound $\mathrm{SLB}(D^{\star})$ corresponds to the moment when the hypotheses $f_{k}$ completely separate from each other (see Theorem 1 and 3 in [59]).